# Inferring Piece Value in Chess and Chess Variants

### v1

@@ Abstract
We use logistic regression to estimate the value of the pieces in standard chess and several chess variants, namely Chess 960, Atomic chess, Antichess, and Horde chess. We perform our regressions on several years of data from Lichess, the free and open-source internet chess server. [1] We use the published player ratings to control for the confounding effect of differential player skill. We adjust for the attenuation bias in regressions due to the noise in observed ratings. We find that major piece values, relative to the value of a pawn, are fairly consistent with historical valuation systems. However we find slightly higher value to bishops than knights. We find that piece values are smaller, in absolute value, in Atomic and Antichess than standard chess. We also present approximate values of the pieces to equalize odds when players of varying skill face off.
@@ 1 Introduction
This valuation system is only one of many that have been devised, however, some dating back to the nineteenth century. [46, 21, 5] Maizelis, for example, lists the “relative strengths of the pieces” as 3.5, 3.5, 5, and 10 (or 9.5), values which are determined by their “mobility in open space”. [26] It is noted, however that piece values are relative to their position on the board, and piece values may change in the endgame. Lasker is similarly elusive, giving more than one valuation system in his book. [24, pages 35, 107] It is a simple exercise to compute the average number of squares each piece could reach on an empty board, namely knight: 5.25, bishop: 8.75, rook: 14, queen: 22.75, and king: 6.5625. Under this calculation, which was known in the nineteenth century, the pawn is assigned value 1.75. [48]
@@ 2.2 Atomic chess
Atomic chess is another popular variant on Lichess. [33, page 42] The game is played with the same board, pieces, piece movements, and opening position as standard chess. As in standard chess the goal is to checkmate your opponent, or otherwise remove your opponent’s king from the board via an “explosion”. These are due to the major change in rules from standard chess:
@@ 2.4 Horde
Horde chess is a variant with an asymmetrical starting position. [8][33, page 90] White starts with 36 pawns arranged as in the following board, while black uses the full complement of pieces in their usual arrangement. White wins by checkmating, black wins by eliminating all of white’s pieces. The white pawns on the first or second rank may move one or two squares for their first move. Black may castle under the usual rules.
@@ 5 Summary and Future Directions
    One can pit different chess programs against each other from selected positions with different material imbalances, then perform the regression procedure to check on the material values found here. This would allow one to check the equalizer values, since the opening position with a few piece deletions is likely to be quiescent. It would also be interesting to see if valuations are similar for computer and human players.
@@ Appendix B Regression Calibration with a Single Noisy Covariate
Here we consider the regression calibration method when a single covariate is observed with noise. Suppose that vector $\boldsymbol{x}$ represents an true covariate which we wish to use in a regression. However we instead observe $\hat{\boldsymbol{x}}$, where $x_{1}$ has some added zero-mean noise with standard deviation $\sigma$. One way to estimate the true regression effects against $\boldsymbol{x}$ is to translate the results from a regression against $\hat{\boldsymbol{x}}$.

### v2

@@ Abstract
We use logistic regression to estimate the value of the pieces in standard chess and several chess variants, namely Chess 960, Atomic chess, Antichess, and Horde chess. We perform our regressions on several years of data from Lichess, the free and open-source internet chess server. [1] We use the published player ratings to control for the confounding effect of differential player skill. We adjust for the attenuation bias in regressions due to the noise in observed ratings. We find that major piece values, relative to the value of a pawn, are fairly consistent with historical valuation systems. However we find slightly higher value to bishops than knights. We find that piece values are smaller, in absolute value, in Atomic and Antichess than standard chess. We also present approximate values of the pieces to equalize odds when players of varying skill face off. We briefly consider self-play experiments using the Stockfish engine, which give a contrasting view of piece value.
@@ 1 Introduction
This valuation system is only one of many that have been devised, however, some dating back to the nineteenth century. [47, 22, 6] Maizelis, for example, lists the “relative strengths of the pieces” as 3.5, 3.5, 5, and 10 (or 9.5), values which are determined by their “mobility in open space”. [27] It is noted, however that piece values are relative to their position on the board, and piece values may change in the endgame. Lasker is similarly elusive, giving more than one valuation system in his book. [25, pages 35, 107] It is a simple exercise to compute the average number of squares each piece could reach on an empty board, namely knight: 5.25, bishop: 8.75, rook: 14, queen: 22.75, and king: 6.5625. Under this calculation, which was known in the nineteenth century, the pawn is assigned value 1.75. [49]
@@ 2.2 Atomic chess
Atomic chess is another popular variant on Lichess. [34, page 42] The game is played with the same board, pieces, piece movements, and opening position as standard chess. As in standard chess the goal is to checkmate your opponent, or otherwise remove your opponent’s king from the board via an “explosion”. These are due to the major change in rules from standard chess:
@@ 2.4 Horde
Horde chess is a variant with an asymmetrical starting position. [9][34, page 90] White starts with 36 pawns arranged as in the following board, while black uses the full complement of pieces in their usual arrangement. White wins by checkmating, black wins by eliminating all of white’s pieces. The white pawns on the first or second rank may move one or two squares for their first move. Black may castle under the usual rules.
@@ 4.5 Horde
5 Inferred Value from Self-Play
Above we considered piece value by observing games between human players on Lichess. Here we briefly examine material value as inferred from self-play of the Stockfish chess engine. [2] We set up experiments pitting two instances of Stockfish version 17.1 against each other, starting from modified opening positions. We start from the standard chess opening and remove one or more pieces from one side or another, then observe tens to thousands of matches from that starting position. We performed a number of simulated matches from the standard opening without pieces removed, as well as from each of the 960 possible Chess 960 openings. We experimented using different target Elo settings for the realizations of Stockfish, which we controlled for in our analysis. We remove one of each of the pieces from each side, simulating hundreds of games. We also simulated games were two major pieces are removed from one side or the other.
We nominally set the “Elo rating” of the stockfish engines to one of four values: 1600, 2000, 2400, 2800. This is controlled via the UCI_Elo configuration option. We note, however, that the ratings are calibrated to a certain time control and we did not observe the nominal spread in ratings, though we do see skill monotonic in the nominal ratings. We limit play to 5 seconds per move and a total depth of 7, which is likely the limiting factor.
We perform logistic regression on the outcomes, again adjusting estimated coefficients so they are in Elo-like units. When considering matches with different nominal strength of the two instances of Stockfish, we include terms for engine strength in the regressions. For these we assign a 1 to factor variables for whichever setting was used for white minus that used for black. That is we have four variables, Elo1600 through Elo2800. We initialize these four as zero, then increment the variable by one if that is the setting used for white, and decrement the variable corresponding to black. Because these four variables will always sum to zero, they are not linearly independent, and we only include Elo2000 through Elo2800 in the regressions, keeping Elo1600 as the implicit “base level”.
In Table 15 we report the results of three regressions:
  1. 1.
    In the first regression we consider only the standard opening, with no pieces removed. We include terms only for three of the four UCI_Elo settings, plus a term for white’s first player advantage. Quoted in Elo-like units, we see a spread between the settings far less than the nominal 400 point spread.
  2. 2.
    In the second regression we look at all standard chess matches, with and without pieces removed, but only consider matches where both engines used the same nominal strength rating. The regression only includes terms for white’s advantage, and the piece imbalance.
  3. 3.
    In the third regression we consider all the standard chess matches we simulated. The regression includes terms for the engine settings, as in the first regression, white’s advantage, and the piece imbalance.
The engine fit values are largely consistent between the first and third regression, and the inferred piece values are largely consistent between the second and thir regressions. However, to our surprise, and consternation, the inferred piece values we fit here on self-play data are quite a bit higher than the “equalizer values” we quoted above based on Lichess data. Possible sources of this discrepancy are: selecting non-quiescent positions in the Lichess data study; more generally some other selection mechanism that biases our Lichess data to less decisive early material imbalances; Stockfish’s greater ability to capitalize on material differences. We leave this question for future research.
Another surprising outcome from these regressions is the rather compressed value of the pieces: a knight and a bishop are worth approximately 3 pawns (2.7 and 2.9, respectively), while a rook is only worth around 3.6 pawns, and queen only around 4.9 pawns. It is not clear why this is so compressed compared to the findings on the Lichess data or the classical piece value heuristics.
Finally we performed regressions on the full set of Standard games where we included a one-hot encoded variable for each square holding a non-king piece in the standard opening. This gives an inferred value for each individual piece. As above we convert the inferred value to Elo-like units. We plot these in Figure 13, showing the piece value at the corresponding square. The regression is on white’s probability of winning, but we flip the sign for black’s pieces so they express their value to black’s chances of winning. We observe the following:
  1. 1.
    The king and queen pawns are worth more than the knight and bishop pawns. Not surprisingly the rook pawns worth less than the others, since they can only attack one other pawn. Black’s rook pawns are worth more than white’s, perhaps for their defensive value.
  2. 2.
    We see greater value to bishops than knights. However, there is greater value to queen side knights and bishops than king’s side. The queen side knight is worth more than the king side bishop.
  3. 3.
    White’s pieces have a higher fit value than black’s pieces. Even though we control for white’s first player advantage, this finding likely reflects the sensitivity of white’s first player advantage to material imbalances.
Table 15: Inferred value of engine strength and material imbalance is shown via three regressions on Stockfish self-play data for standard chess. Standard errors are given in parentheses.
Dependent variable:
Game Outcome
All Pieces Piece Ablations Piece Ablations
(1) (2) (3)
White Advantage 16 (1) 14 (1) 14 (1)
Nominal Elo 2000 215 (3) 251 (2)
Nominal Elo 2400 501 (3) 520 (2)
Nominal Elo 2800 657 (4) 658 (3)
$\Delta$ Pawn 85 (3) 93 (2)
$\Delta$ Knight 245 (2) 250 (2)
$\Delta$ Bishop 265 (3) 274 (2)
$\Delta$ Rook 321 (3) 333 (2)
$\Delta$ Queen 455 (7) 460 (4)
Simulated Matches 171,000 180,400 461,000
Figure 13: The inferred value of each piece in the standard opening are shown. Regressions on Stockfish self-play data, including all 461,000 games.
We also performed self-play experiments on each of the Chess 960 openings, with and without pieces removed. In Table 16 we report the results of three regressions, as we did for standard chess:
  1. 1.
    The first regression considers all Chess 960 openings, with no pieces removed, including terms only for the UCI_Elo settings and white’s first player advantage. Again the spread between settings is larger than nominal.
  2. 2.
    The second regression considers all Chess 960 openings, but looks only at those with equal engine settings, and has terms only for piece imbalance and white’s first player advantage.
  3. 3.
    The third regression looks at all simulated matches, with terms for engine setting, piece imbalance and white’s advantage.
As for the simulations on standard chess, we see that the fit values are consistent between the first and third regression and the second and third regressions. The fits for Chess 960 are very close to those fit from standard chess in Table 15. Again we are surprised to find smaller value to the rook and queen, in pawn terms. We find a knight to be worth approximately 2.6 pawns, a bishop is worth 3.1, a rook is worth only around 3.5 pawns, and a queen only around 5.6 pawns. Here the value of a rook is very depressed compared to classical valuation systems, worth almost as little as a bishop. It is not clear if this is due to decreased mobility of rooks in some large proportion of Chess 960 openings.
In Figure 14 we repeat the plot of Figure 8, with the estimated piece values from Lichess data for standard chess and Chess 960. We add to this plot the estimated piece values for ply zero from the regressions on Stockfish self-play data. We add a dotted line between the Lichess and Stockfish fit values. Other than the value fit for the queen in Chess 960, the self-play fits all represent a sharp departure from the trend seen in the Lichess fit data. Arguably the self-play fit values are too high, or the Lichess fit values are too low for early ply ranged fits. Further research is required to resolve these inconsistencies.
Table 16: Inferred value of engine strength and material imbalance is shown via three regressions on Stockfish self-play data for Chess 960. Standard errors are given in parentheses.
Dependent variable:
Game Outcome
All Pieces Piece Ablations Piece Ablations
(1) (2) (3)
White Advantage 15 (1) 13 (0) 13 (0)
Nominal Elo 2000 247 (3) 268 (1)
Nominal Elo 2400 528 (5) 530 (1)
Nominal Elo 2800 688 (5) 664 (1)
$\Delta$ Pawn 81 (0) 86 (0)
$\Delta$ Knight 222 (1) 227 (0)
$\Delta$ Bishop 262 (1) 267 (0)
$\Delta$ Rook 290 (1) 297 (1)
$\Delta$ Queen 479 (3) 486 (1)
Simulated Matches 96,000 1,958,400 4,896,000
Figure 14: The plot of Figure 8 is augmented with values for zero ply inferred from Stockfish self play experiments.
@@ 6 Summary and Future Directions
    Our experiments with self-play raised more questions than they answered. Further investigation is warranted to see if piece valuations are simply different among human and engine play, or there is some other mechanism that explains the discrepancies we saw. It would be worthwhile to consider self-play experiments on positions that arise in the Lichess data, to compare human and machine performance.
@@ Appendix B Regression Calibration with a Single Noisy Covariate
Here we consider the regression calibration method when a single covariate is observed with noise. Suppose that vector $\boldsymbol{x}$ represents a true covariate which we wish to use in a regression. However we instead observe $\hat{\boldsymbol{x}}$, where $x_{1}$ has some added zero-mean noise with standard deviation $\sigma$. One way to estimate the true regression effects against $\boldsymbol{x}$ is to translate the results from a regression against $\hat{\boldsymbol{x}}$.